# Subsystem localization

### v1

@@ I Introduction
Anderson localization is a key quantum effect where uncorrelated disorder localizes wavefunctions and blocks transport of non-interacting electrons [1, 2, 3]. It creates a mobility edge in three dimensions that separates localized and extended states. A metal-insulator transition occurs when the Fermi energy crosses this edge. In one dimension (1D), all states become localized with any amount of disorder. There is no mobility edge and no transition. Introducing interactions in such disordered Anderson localized systems can stabilize a many-body localized (MBL) phase, where systems fail to thermalize and retain memory of initial conditions [4, 5, 6, 7, 8, 9, 10]. This behavior contradicts the Eigenstate Thermalization Hypothesis (ETH), which posits that isolated non-integrable systems should relax to thermal equilibrium [11, 12, 13, 14]. However, recent studies highlight the avalanche instability within the MBL phase. This instability suggests that rare regions of the system can thermalize and spread thermalization through the system, potentially destabilizing the MBL phase [15, 16, 17].
@@
On the other hand, the presence of correlated disorder can significantly change the Anderson localization scenario in 1D. One can find the localization-delocalization transition and even the existence of the mobility edge, which is also feasible in such systems. One of the prime examples of such systems is the Aubry-André (AA) Hamiltonian, where the uncorrelated ”true” disorder is replaced by a quasi-periodic potential  [18]. One can undergo a delocalization-localization transition by tuning the strength of the quasi-periodic potential. The AA Hamiltonian also has the unique property of self-duality at the critical point, with identical real space and momentum space representations, thus it contains an energy-independent localization transition for the whole spectrum with no mobility edges. It has been shown that by adding different types of perturbations to the AA model, the fine-tuned AA duality conditions can be broken, and an energy-dependent self-duality relation can be established, automatically implying the existence of the mobility edges [19, 20, 21, 22, 23, 24, 25, 26]. With the remarkable progress in cold-atom and ion-trap experiments over the past decade, many of these models have been experimentally realized [27, 28, 29, 30], providing strong motivation to study quasi-periodic systems beyond purely theoretical interest. In this context, the emergence of a non-ergodic metallic phase has been proposed in models with single-particle mobility edges when interactions are introduced. [31, 32, 33, 34, 35]. One of the questions in the field of localization has persisted for some time now, i.e., what happens if one couples localized and extended systems together? In the case of a many-body interacting system, if the degrees of freedom of the extended system are large enough compared to the localized system, the extended system is expected to thermalize the localized one. On the other hand, if the delocalized degrees of freedom are comparable to the localized ones, the outcome is not yet clear. Many efforts have been made to address this question in the last several years [36, 37, 38, 39, 40]. Numerical challenges to tackle the exponentially growing Hilbert space dimension with system size and limitations of analytical tools to solve interacting systems have made this problem extremely challenging. This manuscript has tried to address a similar question in the non-interacting context.
@@ II.1 Model system
Figure 3: (a)-(c): $\nu$ vs. $V$ plots for $t^{\prime}=$ 1, 5 and 10, respectively. The calculations are performed for the total system size (2N) = 1000.
@@ III Static properties
Figure 4: Plots of $IPR_{A}$ as function of $V$ for all $\epsilon$-significant states ($\epsilon$ = 0.3, 0.5 and 0.7). Calculations are performed for the total system size (2N) = 1000 and for $t^{\prime}$ = 1, 5 and 10.
@@ VI Conclusion
In summary, in this work, we study a two-leg ladder model, where one leg is described by an AA-type Hamiltonian (which can undergo a localization-delocalization transition by tuning the strength of the incommensurate potential) and the other leg by the usual nearest-neighbor tight-binding model, which we refer to as the subsystem. By introducing the coupling $t^{\prime}$ between these two legs, we obtain a rich phase diagram as shown schematically in Fig. 2. We notice that beyond a critical coupling strength $t^{\prime}_{c}$, there exists a regime where the subsystem can be completely localized. Also, we find different parameter regimes where the transport can be ballistic, subdiffusive, and superdiffusive in nature. We further explain some of these results using a toy model and also by mapping it to a well-studied 1D GAA model.
@@ Appendix C Understanding Phases for $t^{\prime}<t^{\prime}_{c}$
Figure 15: Variation of $IPR_{A}$ with $V$ for all $\epsilon$-significant states. Calculations are performed for the total system size (2N) = 1000, $t^{\prime}=3$ and $\epsilon=0.5$

### v2

@@ I Introduction
Anderson localization is a key quantum effect where uncorrelated disorder localizes wavefunctions and blocks transport of non-interacting electrons [1, 2, 3]. It creates a mobility edge in three dimensions that separates localized and extended states. A metal-insulator transition occurs when the Fermi energy crosses this edge. In one dimension (1D), all states become localized with any amount of disorder. There is no mobility edge and no transition. Introducing interactions in such disordered Anderson localized systems can stabilize a many-body localized (MBL) phase, where systems fail to thermalize and retain memory of initial conditions [4, 5, 6, 7, 8, 9, 10, 11]. This behavior contradicts the Eigenstate Thermalization Hypothesis (ETH), which posits that isolated non-integrable systems should relax to thermal equilibrium [12, 13, 14, 15]. However, recent studies highlight the avalanche instability within the MBL phase. This instability suggests that rare regions of the system can thermalize and spread thermalization through the system, potentially destabilizing the MBL phase [16, 17, 18].
@@
On the other hand, the presence of correlated disorder can significantly change the Anderson localization scenario in 1D. One can find the localization-delocalization transition and even the existence of the mobility edge, which is also feasible in such systems. One of the prime examples of such systems is the Aubry-André (AA) Hamiltonian, where the uncorrelated ”true” disorder is replaced by a quasi-periodic potential  [19]. One can induce a delocalization–localization transition by tuning the strength of the quasi-periodic potential, a phenomenon that holds significance in various physical contexts  [20, 21, 22, 23, 24]. The AA Hamiltonian also has the unique property of self-duality at the critical point, with identical real space and momentum space representations, thus it contains an energy-independent localization transition for the whole spectrum with no mobility edges. It has been shown that by adding different types of perturbations to the AA model, the fine-tuned AA duality conditions can be broken, and an energy-dependent self-duality relation can be established, automatically implying the existence of the mobility edges [25, 26, 27, 28, 29, 30, 31, 32]. With the remarkable progress in cold-atom and ion-trap experiments over the past decade, many of these models have been experimentally realized [33, 34, 35, 36], providing strong motivation to study quasi-periodic systems beyond purely theoretical interest. In this context, the emergence of a non-ergodic metallic phase has been proposed in models with single-particle mobility edges when interactions are introduced. [37, 38, 39, 40, 41]. One of the questions in the field of localization has persisted for some time now, i.e., what happens if one couples localized and extended systems together? In the case of a many-body interacting system, if the degrees of freedom of the extended system are large enough compared to the localized system, the extended system is expected to thermalize the localized one. On the other hand, if the delocalized degrees of freedom are comparable to the localized ones, the outcome is not yet clear. Many efforts have been made to address this question in the last several years [42, 43, 44, 45, 46]. Numerical challenges to tackle the exponentially growing Hilbert space dimension with system size and limitations of analytical tools to solve interacting systems have made this problem extremely challenging. This manuscript has tried to address a similar question in the non-interacting context.
@@ II.1 Model system
Figure 3: (a)-(c): $\nu$ vs. $V$ plots for $t^{\prime}=$ 1, 5 and 10, respectively. The calculations are performed for the total subsystem size N = 500.
@@ III Static properties
Figure 4: Plots of $IPR_{A}$ as function of $V$ for all $\epsilon$-significant states ($\epsilon$ = 0.3, 0.5 and 0.7). Calculations are performed for the total subsystem size N = 500 and for $t^{\prime}$ = 1, 5 and 10.
@@ VI Conclusion
In summary, in this work, we study a two-leg ladder model, where one leg (called bath) is described by an AA-type Hamiltonian (which can undergo a localization-delocalization transition by tuning the strength of the incommensurate potential) and the other leg by the usual nearest-neighbor tight-binding model, which we refer to as the subsystem. By introducing the coupling $t^{\prime}$ between these two legs, we obtain a rich phase diagram as shown schematically in Fig. 2. We notice that beyond a critical coupling strength $t^{\prime}_{c}$, there exists a regime where the subsystem can be completely localized. Also, we find different parameter regimes where the transport can be ballistic, subdiffusive, and superdiffusive in nature. However, it is important to note that there is an intermediate region, labeled as P4 phase in Fig. 2, where the subsystem exhibits a complex phase structure. Here it may show super- or subdiffusive behavior based on parameter values and time scale of investigation. A more detailed study is required to gain better insight into the nature of phases in this crossover region. We further explain some of these results using a toy model and also by mapping it to a well-studied 1D GAA model.
@@ Appendix C Understanding Phases for $t^{\prime}<t^{\prime}_{c}$
Figure 15: Variation of $IPR_{A}$ with $V$ for all $\epsilon$-significant states. Calculations are performed for the total subsystem size N = 500, $t^{\prime}=3$ and $\epsilon=0.5$